# Tree Search for Language Model Agents

### v1

@@ 5.1 Ablations
Figure 3: Success rate (SR) and relative change over the baseline ($\Delta$) on a subset of 200 VWA tasks with varying search depth ($d$) and branching factor ($b$). $d=0$ indicates no search is performed. All methods use a maximum search budget $c=20$.
@@
Trajectory-level reranking
An alternative to tree search would be to generate multiple trajectories, re-rank, and commit to the best one as scored by the value function, similar to the methods proposed in Chen et al. (2024b) and Pan et al. (2024) without their Reflexion (Shinn et al., 2024) component. This is a less practical method, as it is harder to prevent destructive actions from being executed (see Sec. 5.4 for more discussion) as the agent is required to take the trajectory to completion before it can be evaluated. It is also a more limited form of search, as it only considers entire trajectories and cannot backtrack to prune bad branches. Nevertheless, we perform an ablation where we sample 3 trajectories from the GPT-4o agent (with nucleus sampling (Holtzman et al., 2020) at each step using a temperature of 1.0 and top-$p$ of 0.95) and use the same value function to re-rank the trajectories, picking the best one out of 3. We choose 3 as this is the same number of trajectories that Pan et al. (2024) use (performance for their approach peaks at 3 trajectories). This re-ranking baseline achieves a success rate of 28.5%, which is worse than our approach with search budget $c\geq 5$ (Fig. 3). It is also substantially worse than our approach with $c=20$, which achieves a success rate of 37.0% on the ablation subset, a relative 29.8% increase over re-ranking.
@@ 5.2 Success Rate Breakdown
The VWA benchmark includes labels for the action difficulty of each task. These labels are human annotated, and roughly indicate the number of actions a human would need to take to solve the tasks: easy tasks require 3 or fewer actions, medium tasks require 4–9 actions, and hard tasks demand 10 or more. These guidelines are approximate and devised by the human annotators of VWA, so there may exist more optimal solutions in practice. The increase in success rate from introducing search is summarized in Tab. 3. Introducing search improves performance across all difficulty levels, but it has a much larger gain on tasks of medium action difficulty, with a relative increase of 75% in success rate (from 12.7% to 22.2%). We hypothesize that this is because our search parameters (max depth $d=5$) are beneficial for a large proportion of medium difficulty tasks. Conversely, achieving even better performance on hard tasks may require search over deeper trees. Easy tasks likely do not benefit as much from search, as they generally involve less multi-step planning (some can be solved with 1 or 2 actions), and baselines already have higher success rates.
@@
Tables 5 and 5 summarize the success rates across the various websites in the VWA and WA benchmarks. We observe an improvement in success rates across the board, demonstrating that our method generalizes across sites. Specifically, the increase is most substantial on the Classifieds and Shopping sites in VWA, with relative increases of 44% and 45%, respectively. Similarly, the CMS site in the WA benchmark shows a significant relative improvement of 50%.
@@ 5.3 Qualitative Results
In this section, we show several qualitative examples of agent trajectories, and identify various failure modes that can be solved through incorporating search.
@@
Many tasks in VWA and WA require an agent to keep a persistent memory of multiple previous actions and observations. A common failure mode amongst agents without search is that they tend to undo previous actions, or get stuck in loops (see Appendix C.4 of Koh et al. 2024). An example for VWA shopping task #256 is shown in Fig. 1, where the agent is tasked to add two different types of canned fruit from the same brand to the comparison list. The baseline agent successfully adds the first item, but fails to navigate to the second item, as it returns to the homepage in step 3 and gets confused. This is an example of compounding error leading to overall task failure, which is fairly common in existing baseline agents.
When search is introduced, the agent is now able to explore other plausible trajectories and backtrack when those eventually result in failure. The same GPT-4o agent with search is now able to find a successful multi-step trajectory for this task, which involves adding the first item (action #1 in Fig. 1), typing in a search query (action #6), and navigating and adding the correct second item to the comparison list (action #8).
@@
An inherent issue with sampling actions from language models is that we are sampling from a distribution over text, and the first sample we generate may not always be the best action to take in the environment. Search allows us to evaluate each generated action concretely by executing it in the simulator, and use the received environmental feedback to make better decisions. One example is VWA classifieds task #48 (Fig. 4), which is to find a post containing a particular image. If the agent executed the first sampled action at each step, it would lead to a failure. Search allows the agent to enumerate all possibilities after executing the action and receiving environment feedback, which allows it to select the best (and in this case, successful) trajectory.
@@ 5.4 Limitations
While we have shown that introducing search to LM agents obtains promising results on web tasks, it does come with some practical limitations. In this section we discuss some common failure modes and possible ways to address these.
@@
Introducing search allows us to expend more compute at inference time to extract stronger results from the pretrained LM agent. However, this results in trajectories taking significantly longer to execute, as the agent has to perform more exploration and hence more inference calls to the LM. For example, a search budget of $c=20$ implies that an agent with search could potentially expand up to 20 states in each search iteration, which would use up to 20×20$\times$20 × more LM calls than an agent without search. Research on improving the efficiency and throughput of machine learning systems (Leviathan et al., 2023; Dao et al., 2022; Dao, 2023) will likely help with optimizing this, but for practical deployment one may need to carefully set the search parameters $b$, $d$, and $c$ to balance between achieving better results and overall time spent completing a task.
@@
For real world deployment, we will need to restrict the search space to actions that are not destructive. Destructive actions are defined as actions that will irreversibly change the state of the website and are difficult to backtrack from. For example, placing an order on an e-commerce site is typically very difficult to automatically undo.
One way to address this is to introduce a classifier that predicts when certain actions are destructive, and prevent node expansion for those states. If we have specific domain knowledge about the downstream application (e.g., we know certain pages should be off limits), such rules can be manually enforced with high accuracy. One advantage of our method over trajectory level reranking (Sec. 5.1) is that it is easier to incorporate such a constraint: it can be directly integrated into the value function to prevent execution of dangerous actions. Another direction to handle this would be to train a world model (Ha & Schmidhuber, 2018) that we can use for simulations, and search over this rather than explore the real world. Search may also be more easily implemented in offline settings where actions are non-destructive as they can always be undone or reset, such as programming (Jimenez et al., 2023; Yang et al., 2024) or Microsoft Excel (Li et al., 2024).
@@ 6 Conclusion
In this paper, we introduced an inference-time search algorithm designed to enhance the capabilities of language model agents on realistic web tasks. Our approach integrates best-first tree search with LM agents, enabling them to explore and evaluate multiple action trajectories to achieve superior performance on web tasks. This is the first time search has shown to significantly improve the success rates of LM agents on realistic web environments, as demonstrated on the (Visual)WebArena benchmarks. Our search procedure is general, and it will be valuable to apply it to other domains in future work. We believe that inference-time search will be a key component for building capable agents that can plan, reason, and act autonomously to perform computer tasks.
@@ Acknowledgments
We thank Andy Liu, Zhiruo Wang, Alex Xie, Shuyan Zhou, and many others for helpful feedback and discussions on previous versions of this paper. We thank Wendy Kua for helping with the figures.
@@ A.1 Qualitative Examples
A.2 Implementation Details
A.2.1 Language Model Agents
@@
For all experiments, we use a webpage viewport width of 1280, a viewport height of 2048, and truncate text observations to 3840 tokens. We sample from models using nucleus sampling (Holtzman et al., 2020) with a temperature of 1.0 and a temperature of 1.0 and a top-p of 0.95. The system message used in all our experiments is provided in Fig. 7. This instructs the agent with the guidelines for the web navigation task, and list out all the possible actions that it can perform.
@@ A.3 Search Algorithm
6:     $s_{p}\leftarrow\text{pop}(\mathcal{F})$
@@
24:           Add new candidate state to the frontier: $\mathcal{F}\leftarrow\mathcal{F}\cup s_{p}^{i}$

### v2

@@ 5.1 Ablations
Table 3: Success rate (SR) and relative change ($\Delta$) over the baseline without search on a subset of 200 VWA tasks with varying search depth ($d$) and branching factor ($b$). $d=0$ indicates no search is performed. All methods use a max search budget $c=20$.
@@
Value Function SR ($\uparrow$)
None (no search) 24.5%
LLaVA (w/ SC, $n=20$) 30.0%
GPT-4o (no SC) 28.5%
GPT-4o (w/ SC, $n=5$) 32.5%
GPT-4o (w/ SC, $n=20$) 37.0%
Groundtruth 43.5%
Table 4: Success rate of the GPT-4o agent with different value functions.
Varying the value function
We ablate the multimodal model used for the value function, swapping out GPT-4o for (1) the LLaVA-v1.6-34B (Liu et al., 2024a) multimodal model prompted zero-shot (with just the current observation) and (2) the groundtruth reward from VWA (which is a sparse reward signal that returns either 0 or 1), and (3) GPT-4o without self-consistency. The results are summarized in Tab. 4. We find that the GPT-4o value function significant outperforms the LLaVA model, improving the result of the agent from 30.0% to 37.0%. The groundtruth reward function achieves a success rate of 43.5%. These results suggest that there is still significant headroom in improving the search algorithm by training better value functions. We also observe that self-consistency is essential for the performance of search ($28.5\%\rightarrow 37.0\%$), which we attribute to it enabling marginalization over multiple reasoning chains, reducing noise during state evaluation.
@@ 5.2 Success Rate Breakdown
The VWA benchmark includes labels for the action difficulty of each task. These labels are human annotated, and roughly indicate the number of actions a human would need to take to solve the tasks: easy tasks require 3 or fewer actions, medium tasks require 4–9 actions, and hard tasks demand 10 or more. These guidelines are approximate and devised by the human annotators of VWA, so there may exist more optimal solutions in practice. The increase in success rate from introducing search is summarized in Tab. 5. Introducing search improves performance across all difficulty levels, but it introduces much larger gains on tasks of medium action difficulty, with a relative increase of 75% in success rate (from 12.7% to 22.2%). We hypothesize that this is because our search parameters (max depth $d=5$) are beneficial for a large proportion of medium difficulty tasks. Conversely, achieving even better performance on hard tasks may require search over deeper trees. Easy tasks likely do not benefit as much from search, as they generally involve less multi-step planning (some can be solved with 1 or 2 actions), and baselines already have higher success rates.
@@
Tables 7 and 7 summarize the success rates across the various websites in the VWA and WA benchmarks. We observe an improvement in success rates across the board, demonstrating that our method generalizes across sites. Specifically, the increase is most substantial on the Classifieds and Shopping sites in VWA, with relative increases of 44% and 45%, respectively. Similarly, the CMS site in the WA benchmark shows a significant relative improvement of 50%.
@@ 5.3 Qualitative Results
In this section, we discuss some qualitative examples of agent trajectories, and identify various failure modes that are solved when incorporating search.
@@
Many tasks in VWA and WA require an agent to keep a persistent memory of multiple previous actions and observations. A common failure mode amongst agents without search is that they tend to undo previous actions, or get stuck in loops (see Appendix C.4 of Koh et al. 2024). An example for VWA shopping task #256 is shown in Fig. 1, where the agent is tasked to add two different types of canned fruit from the same brand to the comparison list. The baseline agent successfully adds the first item, but fails to navigate to the second item, as it returns to the homepage in step 3 and gets confused. This is an example of compounding error leading to overall task failure, which is fairly common in existing baseline agents without search.
When search is introduced, the agent explores other plausible trajectories and backtrack when those eventually result in failure: the same GPT-4o agent with search is able to find a successful multi-step trajectory for the same task, which involves adding the first item (action #1 in Fig. 1), typing in a search query (action #6), and adding the correct second item to the comparison list (action #9).
@@
An inherent issue with sampling actions from language models is that we are sampling from a distribution over text, and the first sample we generate may not always be the best action to take in the environment. Search allows us to evaluate each generated action concretely by executing it in the simulator, and use the received environmental feedback to make better decisions. One example is VWA classifieds task #48 (Fig. 3), which is to find a post containing a particular image. If the agent executes the first sampled action at every step (i.e., the sequence in the top row), it results in failure. Search allows the agent to enumerate all possibilities by executing plausible actions and receiving environment feedback, which allows it to select the best (and in this case, successful) trajectory.
@@ 5.4 Limitations
While we have shown that introducing search to language model agents achieves promising results on web tasks, it does come with some practical limitations:
@@
Introducing search allows us to expend more compute at inference time to extract stronger results from the baseline LM agent. However, this results in trajectories taking significantly longer to execute, as the agent has to perform more exploration and hence more inference calls to the LM. For example, a search budget of $c=20$ implies that an agent with search could potentially expand up to 20 states in each search iteration, which would use up to 20×20$\times$20 × more LM calls than an agent without search. Research on improving the efficiency and throughput of machine learning systems (Leviathan et al., 2023; Dao et al., 2022; Dao, 2023) will likely help with optimizing this, but for practical deployment one may need to carefully set the search parameters $b$, $d$, and $c$ to balance between achieving better results and overall time spent completing a task.
@@
For real world deployment, we will need to restrict the search space to actions that are not destructive. Destructive actions are defined as actions that will irreversibly change the state of the website and are difficult to backtrack from. For example, placing an order on an e-commerce site is typically very difficult to automatically undo. One way to address this is to introduce a classifier that predicts when certain actions are destructive, and prevent node expansion for those states. If we have specific domain knowledge about the downstream application (e.g., we know certain pages should be off limits), such rules can be manually enforced with high accuracy. One advantage of tree search is that it is easier to incorporate such a constraint: it can be directly integrated into the value function to prevent execution of dangerous actions. Another direction to handle this would be to train a world model (Ha & Schmidhuber, 2018) that we can use for simulations during search. Search may also be more easily implemented in offline settings where actions are non-destructive as they can always be undone or reset, such as programming (Jimenez et al., 2023; Yang et al., 2024) or Microsoft Excel (Li et al., 2024).
@@ 6 Conclusion
In this paper, we introduced an inference-time search algorithm designed to enhance the capabilities of language model agents on realistic web tasks. Our approach integrates best-first tree search with LM agents, enabling them to explore and evaluate multiple action trajectories to achieve superior performance on web tasks. This is the first time search has been shown to significantly improve the success rates of LM agents on realistic web environments, as demonstrated on the (Visual)WebArena benchmarks. Our search procedure is general, and it will be valuable to apply it to other domains in future work. We believe that inference-time search will be a key component for building capable agents that can plan, reason, and act autonomously to perform computer tasks.
@@ Acknowledgments
We thank Andy Liu, Saujas Vaduguru, Zhiruo Wang, Alex Xie, Shuyan Zhou, and many others for helpful feedback and discussions on previous versions of this paper, and Santiago Cortes-Gomez for infrastructure support. We thank Wendy Kua for helping with the figures.
Ethics Statement
As an active area of machine learning research, language model web agents present both opportunities and potential ethical considerations. Improved web agents could improve accessibility for users with disabilities, automate repetitive or tedious tasks, and potentially democratize access to complex web platforms. Our search method contributes towards making such benefits more reliable and widely available by improving the robustness and success rate of language model agents. However, we acknowledge several potential ethical considerations:
  • Intended uses. Our work is a research product that aims to advance the development of web agents that can help augment humans by automating computer tasks. It is not in its current state intended for deployment in practical scenarios. However, we acknowledge that as they get better, enhanced web agents might be leveraged for malicious purposes, such as more sophisticated phishing attempts or automated attacks on web services. As with all emerging technologies, developers deploying these technologies should incorporate consider potential misuse scenarios and implement the appropriate safeguards.
  • Economic impact. As web agents become more capable, there may be concerns about job displacement for roles that involve web-based tasks. We believe that web agents will augment human capability, and will be able to improve the overall quality of work by automating tedious computer tasks. However, as this technology starts being deployed more broadly, researchers and developers should proactively consider how to manage this transition and support affected workers.
  • Fairness and bias. As with any modern AI system, web agents may inherit or amplify biases present in their training data or underlying language models. Care must be taken to assess and mitigate unfair treatment or representation of different user groups. As an inference time algorithm, our approach can easily be applied to any off-the-shelf language model, and will likely benefit from upstream efforts on language model safety and alignment.
Our approach also potentially provides a framework that could help address some of these concerns. The value function in our tree search algorithm offers a natural way to encode safety constraints at inference time. For example, classifiers can be integrated with our proposed value function to prevent destructive actions or violations of privacy and security policies. We encourage further research into the ethical implications of web agents, and the development of guidelines and best practices for the responsible deployment of web agents.
Reproducibility Statement
To ensure reproducibility of our results, we provide details of the tree search algorithm within this paper (Sec. 3.3) as well as more formally in Appendix. A.4. The details of the models used are also provided in Sec. 4.1. All other prompts and implementation details necessary to reproduce our results are provided in Appendix A.3. In order to ensure long term reproducibility, we also provide results for a setting with open model weights for longer term reproducibility, the Llama-3 agent with the LLaVA value function (Appendix. A.2.1), in addition to the API-based agents (which at present achieve higher performance than open sourced alternatives). All of our code and experiment launch scripts are open sourced and made publicly available on GitHub at jykoh.com/search-agents.
@@ A.1 Qualitative Examples
A.2 Additional Ablations
A.2.1 Value Function Ablations
Agent Model Value Function Max Steps No Search + Search $\Delta$
VWA Llama-3-70B-Instruct (Koh et al., 2024) - 30 9.8% - -
GPT-4o + SoM (Koh et al., 2024) - 19.8% - -
Llama-3-70B-Instruct + captions LLaVA-1.6-34B 5 7.6% 13.5% +77.6%
Llama-3-70B-Instruct + captions GPT-4o 7.6% 16.7% +119.7%
Llama-3.1-70B-Instruct + captions GPT-4o 9.1% 16.2% +78.0%
GPT-4o-mini + SoM GPT-4o-mini 9.1% 14.4% +58.2%
GPT-4o + SoM GPT-4o 18.9% 26.4% +39.7%
WA GPT-4o (Zhou et al., 2024b) - 30 13.1% - -
GPT-4 + Reflexion (Pan et al., 2024) - 15.6% - -
AutoWebGLM (Lai et al., 2024) - 18.2% - -
AutoEval (Pan et al., 2024) - 20.2% - -
BrowserGym (Drouin et al., 2024) - 23.5% - -
SteP (Sodhi et al., 2024) - 35.8% - -
Llama-3-70B-Instruct GPT-4o 5 7.6% 10.1% +32.3%
GPT-4o GPT-4o 15.0% 19.2% +28.0%
Table 8: Success rates (SR) and relative change ($\Delta$) for baseline models and models that employ search on the VisualWebArena (VWA) (Koh et al., 2024) and WebArena (WA) (Zhou et al., 2024b) benchmarks. We also show other published approaches. Search substantially improves our baseline models, setting a new state-of-the-art on VWA.
In Sec. 4.2 of the main paper, we experimented with using gpt-4o as our value function. In Tab. 8, we present results using different language models as the agent models and the value functions. We observe that our tree search algorithm is effective across a range of different model sizes and capabilities. In particular, our approach applied to the Llama-3-70B-Instruct and LLaVA-1.6-34B value function yields a 77.6% relative improvement over the baseline Llama-3-70B-Instruct agent on VWA (7.6% to 13.5%), and is a fully open sourced and reproducible baseline. For the GPT-4o-mini model (a relatively weaker model compared to GPT-4o) we also observed improvements when it is used as both the agent model and the value function, improving performance by 58.2% over the no-search baseline on VWA (9.1% to 14.4%).
A.2.2 Comparison to Trajectory-Level Reranking
Figure 6: Success rate of a trajectory re-ranking approach compared to our approach.
An alternative to tree search would be to generate multiple trajectories, re-rank, and commit to the best one as scored by the value function, similar to the methods proposed in Chen et al. (2024b) and Pan et al. (2024) without their Reflexion (Shinn et al., 2024) component. This is a less practical method, as it is harder to prevent destructive actions from being executed (see Sec. 5.4 for more discussion) as the agent is required to take the trajectory to completion before it can be evaluated. It is also a more limited form of search, as it only considers entire trajectories and cannot backtrack to prune bad branches. Nevertheless, we perform an ablation where we sample $n$ trajectories from the GPT-4o agent (with nucleus sampling (Holtzman et al., 2020) at each step using a temperature of 1.0 and top-$p$ of 0.95) and use the same value function to re-rank the trajectories, picking the best one out of $n$.
We observe that this re-ranking baseline starts to plateau around 7 runs, which achieves a success rate of 30%. This underperforms our approach with search budget $c\geq 5$ (Fig. 2). It is also substantially worse than our approach with $c=20$, which achieves a success rate of 37.0% on the ablation subset.
A.3 Implementation Details
A.3.1 Language Model Agents
@@
For all experiments, we use a webpage viewport width of 1280, a viewport height of 2048, and truncate text observations to 3840 tokens. We sample from models using nucleus sampling with a temperature of 1.0 and a temperature of 1.0 and a top-p of 0.95. The system message used in all our experiments is provided in Fig. 7. This instructs the agent with the guidelines for the web navigation task, and list out all the possible actions that it can perform.
@@ A.4 Search Algorithm
6:     $s_{p},v_{\text{prev}}\leftarrow\text{pop}(\mathcal{F})$
@@
24:           Add new candidate state and the current value to the frontier: $\mathcal{F}\leftarrow\mathcal{F}\cup(s_{p}^{i},v_{p})$